### v1

@@ Introduction.
The spectrum calculations can be emulated varying the $\bm{\theta}$ parameter in $H$, as shown in Fig. 3. The sources (fixed) and $\mathrm{Im}\,E_{\alpha}^{tr}$ are the same as in Fig. 2. In the top panels, $E$ and $\lambda_{4}$ are the emulation variables, while $\lambda$ is fixed. Two emulators with different $\lambda$ are trained (see the plot titles). We sample $N_{b}=60$ points in the $\mathrm{Re}\,E_{\alpha}^{tr}$-$\lambda_{4}$ space using Latin hypercube sampling (LHS) Tang (1993) with $\mathrm{Re}\,E_{\alpha}^{tr}\in[-20,50]$ MeV and $\lambda_{4}\in[-0.5,0.5]$. The spectrum emulated at a randomly chosen $\lambda_{4}$ is plotted for each emulator. In the bottom panels, $\lambda$ is included as an emulation variable. Thus, only one emulator is needed and trained, with $N_{b}=60$ training points sampled using LHS in the $\mathrm{Re}\,E_{\alpha}^{tr}$-$\lambda_{4}$-$\lambda$ space with $\lambda\in[-2.15,1]$ ($\mathrm{Re}\,E_{\alpha}^{tr}$ and $\lambda_{4}$ ranges unchanged). The emulator is tested at two parameter sets (see the plot titles). In the figure, the left panels have bound dimers and none for the right. These emulated spectra are qualitatively similar to those in Fig. 2, including the agreement between the emulated physical states and the exact results. These plotted spectra demonstrate the spectrum emulation capability of our method. However, deep into the complex plane (10s MeV below the real axis), the supposedly BCT eigenvalues are scattered somewhat. This is not problematic for emulating real-$E$ observables and near-axis resonances but can make isolating possible broad-resonance poles difficult. This phenomenon needs further studies.
@@
Figure 4: The nonBorn term of the particle-dimer scattering $T$-matrix are emulated by two emulators with different $\mathrm{Im}\,E^{tr}_{\alpha}$. For each curve, $\lambda$ and thus $B_{2}$ are fixed (see the legend), but a sample of $\lambda_{4}$ values is chosen to test the emulators. The means of the sample of relative errors are plotted.
@@
To emulate particle-dimer scattering amplitudes, we must vary the source parameters, including the interaction couplings and the scattering energy $E_{rel}$ between the particle and dimer, in addition to those in $H$. For on-shell scatterings, as computed in the emulation stage, the real-valued $E_{rel}$ is related to $E$ via $E_{rel}-B_{2}=E$. However, the training calculations treat them as separate emulation variables. Thus, we sample $N_{b}=128$ training points using LHS in the $\mathrm{Re}\,E$-$E_{rel}$-$\lambda$-$\lambda_{4}$ space, with $E_{rel}\in[0,60]$ MeV, $\lambda$ in a range in which $B_{2}\in[2,10]$ MeV, and $\mathrm{Re}\,E_{\alpha}^{tr}$ and $\lambda_{4}$ in the same ranges as for Fig. 3.
@@
Two different emulators with $\mathrm{Im}\,E^{tr}_{\alpha}=3$ or $10$ are trained. They are checked against the exact on-shell $T$-matrices at a sample of 1000 points in the $E_{rel}$-$\lambda$-$\lambda_{4}$ space. Figure 4 demonstrates emulation performance by showing the relative errors of emulating the $T$-matrix’s nonBorn term versus $E_{rel}$. Each curve represents the mean of the relative errors from a sub-group of testing points with the same $B_{2}$ (see the legend) but different $\lambda_{4}$. The errors increase towards the particle-dimer threshold, similar to the behaviors of the rational approximations of functions near the branch points. Based on the comparisons, reducing $\mathrm{Im}\,E^{tr}_{\alpha}$ systemically improves the emulation, particularly around the branch points.
@@ Discussion.
Our method has two components: the training calculations performed with complex $E_{\alpha}^{tr}$ and real $\bm{\theta}_{\alpha}^{tr}$, and the emulations that extrapolate the complex-$E$ training results to the other parts of the complex plane and interpolate (or extrapolate) in $\bm{\theta}$. The computing costs for emulation are low, but those of the training calculations increase with system size. Still, the complex-$E$ solutions have been computed in the continuum studies Carbonell et al. (2014) employing the CE Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) and LIT methods Efros (1985); Efros et al. (1994, 2007); Orlandini et al. (2014); Sobczyk et al. (2021, 2024); Bonaiti et al. (2024) for few- and many-body systems. Our method could thus apply to these systems as well.
To infer the real-$E$ results from the complex-$E$ solutions, the CE method uses rational-approximation based extrapolants Schlessinger and Schwartz (1966); Schlessinger (1968a, b); McDonald and Nuttall (1969); Uzu et al. (2003); Deltuva and Fonseca (2012, 2013a, 2013b, 2014) while LIT fits real-$E$ observables via integral transforms Efros (1985); Efros et al. (1994, 2007); Orlandini et al. (2014); Sobczyk et al. (2021, 2024); Bonaiti et al. (2024). In our method, the RBM emulator is used to realize this connection. It creates an NHQM approach that computes not only observables but also spectra in which bound and resonance states are separated from the BCT states. All three methods discussed here thus belong to the type-II methods mentioned earlier—an insight of this work.
Importantly, our method enables emulating the CE and LIT calculations. It provides fast interpolations of $\mathcal{A}$ at general $\bm{\theta}$ and $E$, given that $E$ is close to $E_{\alpha}^{tr}$. The complex-$E$ results at a given $\bm{\theta}$ can then be used as the inputs for the complex-$E$-to-real-$E$ procedures in CE and LIT. We thus create fast interpolations of the real-$E$ observables in the $\bm{\theta}$ space. See Ref. Zhang (2024) for details.
This work not only adds a new tool for studying continuum physics but also expands the MOR literature by applying the RBM method to project an operator with a continuum spectrum into a finite matrix; previous MOR literature Antoulas (2005) mainly focuses on dimension reduction of large matrices. Our study also suggests a connection between this RBM projection and (near-)optimal rational approximation of functions with branch points Nakatsukasa et al. (2018); Trefethen et al. (2021); Trefethen (2023); Nakatsukasa et al. (2023).
@@
We derive essential understandings of the new method based on the training calculations with tiny errors (on the order of $10^{-12}$). Meanwhile, the numeric analytical continuation in $E$ is known to be sensitive to the errors in the data to be extrapolated Trefethen (2023). Such sensitivity will be addressed later. It will require the development of regularization methods, perhaps adapted to the specific many-body methods employed to perform the training calculations. For this purpose, the general behavior of the emulations exposed in this work, such as the distribution of eigenvalues, provides valuable prior knowledge for developing the regularization methods Hicks and Lee (2022).
@@
We apply the RBM method to emulate inhomogenous Schrödinger equations in a combined space of the complex $E$ plane and the parameters inside the Hamiltonian and the sources. It creates an NHQM method for extracting and emulating continuum states based on bound-state-like calculations. However, the method differs from previous NHQM methods in that it constructs the subspace by RBM-based projections, while the previous methods form a many-body basis with direct products of single-particle states. Good emulation performances are demonstrated in two and three-body systems as proof of principles. Moreover, the method can help emulate existing continuum calculations, such as those based on CE and LIT methods. Finally, the potential connection between the RBM’s effectiveness in mapping out the solution manifold and the optimality of the rational approximation should be further explored.
@@ Acknowledgements.
Discussions with Dean Lee, Dick Furnstahl, Nobuo Hinohara, Chong Qi, Simin Wang, Alex Gnech, Bijaya Acharya, and Ante Ravlic are greatly appreciated. I also thank Dick for careful reading of the manuscript. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under the FRIB Theory Alliance Award No. DE-SC0013617 and under the STREAMLINE Collaboration Award No. DE-SC0024586.

### v2

@@ Introduction.
These emulated spectra are qualitatively similar to those in Fig. 2, including the agreement between the emulated physical states and the exact results. The results thus demonstrate the spectrum emulation capability of our method. However, deep into the complex plane (10s MeV below the real axis), the supposedly BCT eigenvalues are scattered somewhat. This is not problematic for emulating real-$E$ observables and near-axis resonances but can make it difficult to separate possible broad resonancs from the nearby states that supposedly represent BCT (s). This phenomenon needs further studies.
@@
Figure 4: The non-Born term of the particle-dimer scattering $T$-matrix are emulated by two emulators with different $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$. For each curve, $\lambda$ and thus $B_{2}$ are fixed (see the legend), but a sample of $\lambda_{4}$ values is chosen to test the emulators. The means of the sample of relative errors are plotted.
@@
To compute and emulate particle-dimer scattering amplitudes, we must vary the source parameters, including the interaction couplings and the scattering energy $E_{\mathrm{rel}}$ between the particle and dimer, in addition to those in $H$. For on-shell scatterings, as computed in the emulation stage, the real-valued $E_{\mathrm{rel}}$ is related to $E$ via $E_{\mathrm{rel}}-B_{2}=E$. However, the training calculations treat them as separate emulation variables. Thus, we sample $N_{b}=128$ training points using LHS in the $(\mathrm{Re}\,E,E_{\mathrm{rel}},\lambda,\lambda_{4})$ space, with $E_{\mathrm{rel}}\in[0,60]$ MeV, $\lambda$ in a range in which $B_{2}\in[2,10]$ MeV, and $\mathrm{Re}\,E_{\alpha}^{\mathrm{tr}}$ and $\lambda_{4}$ in the same ranges as for Fig. 3.
@@
Two different emulators with $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}=3$ and $10$ are trained. They are checked against the exact on-shell $T$-matrix non-Born terms at a sample of 1000 points in the $(E_{\mathrm{rel}},\lambda,\lambda_{4})$ space. In Fig. 4, each curve represents the mean of the relative emulation errors from a sub-group of testing points with the same $B_{2}$ but different $\lambda_{4}$. The tiny errors at energies above thresholds suggest robust emulation performances there. Meanwhile, the (relative) errors increase towards the particle-dimer threshold, similar to the behaviors of the rational approximations of functions near their branch points. However, reducing $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ (from $10$ to $3$ MeV here) improves the emulations around thresholds systemically..
@@ Discussion.
Our emulators are trained using the calculations performed with complex $E_{\alpha}^{\mathrm{tr}}$ and real $\bm{\theta}_{\alpha}^{\mathrm{tr}}$. The cost of individual training calculation increases with system size, but its feasibility for few- and many-body systems have been demonstrated in previous studies [68] employing the CE [70, 71, 72, 73, 74, 75, 76, 77, 78]  [^6] Here, training points typically differ only by their $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$ values. and LIT methods [2, 79, 3, 69, 80, 81, 82]. These works underscore this powerful strategy that exploits advanced bound-state solvers to access continuum physics. The choice then becomes the method for inferring the real-$E$ results from the complex-$E$ solutions.
To do so, the CE method uses continued-fraction based extrapolants [70, 71, 72, 73, 74, 75, 76, 77, 78] while LIT fits real-$E$ observables via integral transforms [2, 79, 3, 69, 80, 81, 82]. In contrast, our method uses the RBM emulator to realize this connection, which keeps the number of training calculations minimal and offers extreme interpolating and extrapolating speed. Interestingly, this work creates an NHQM approach that computes both observables and spectra and obtains an insight that the complex-energy based methods belong to the type-II category mentioned earlier.
Importantly, our $\bm{\theta}$-emulation, indispensable for exploring parameter space, can be adapted to work with the CE and LIT methods. It provides fast interpolations of $\mathcal{A}$ at general $\bm{\theta}$ and complex energies (close to $E_{\alpha}^{\mathrm{tr}}$), which can be used as the inputs for the complex-$E$-to-real-$E$ procedures in CE and LIT, creating fast $\bm{\theta}$-emulations of their continuum calculations. See Ref. [4] for details.
Moreover, this work expands the MOR literature by applying the RBM method to project an operator with a continuous spectrum into a finite matrix; previous MOR literature [89] mainly focuses on dimension reduction of large matrices. Our study also shows great similarities between the complex-$E$ emulation and the ORA for functions with branch points [84, 85, 86, 87].
@@
One direction is to implement greedy algorithms [46, 66] for selecting training points (including with random $\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}$), as suggested by the redundancy ($N_{\mathrm{eff}}\ll N_{b}$) observed above. We then obtain optimal bases with minimal training costs. The algorithms would also provide a more sophisticated regularization than the current SVD truncation, further stabilizing the analytical continuation in the complex-$E$ plane as seen in the context of ORA [84]. The continuation, in particular near thresholds, could be also enhanced by enforcing general analytical behaviors in $\begin{bmatrix}\mathcal{A}\end{bmatrix}$ within, e.g., the framework of Bayesian statistics [90]. Lastly, we can extend the framework to treat systems with strong Coulomb effects by employing Coulomb wave functions [3].
@@
We apply the RBM method to emulate inhomogeneous Schrödinger equations in a combined space of the complex-$E$ plane and other parameters in the equations. It creates an NHQM method for extracting and emulating continuum states based on bound-state-like calculations. However, the method differs from previous NHQM methods in that it constructs the subspace by RBM-based projections, while the previous methods form a many-body basis with direct products of single-particle states. Good emulation performances are demonstrated in two and three-body systems as proof of principles. The method can also help emulate existing continuum calculations, such as those based on CE and LIT methods. Finally, the potential connection between our emulation and ORA should be further explored.
@@ Acknowledgements.
Discussions with Dean Lee, Dick Furnstahl, Nobuo Hinohara, Chong Qi, Simin Wang, Alex Gnech, Bijaya Acharya, and Ante Ravlic are greatly appreciated. I also thank Dick Furnstahl for careful reading of the manuscript. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, under the FRIB Theory Alliance Award No. DE-SC0013617 and under the STREAMLINE Collaboration Award No. DE-SC0024586 (Michigan State University).